1. X: the number of automobile accidents per year in Virginia.
1
Expert's answer
2022-01-31T15:53:36-0500

The question is incomplete, assume it as:


Let the total number of accidents in Virginia are 3500.

And X: the number of automobile accidents per year in Virginia, "n(X)=2800"

and Y: the number of other accidents per year in Virginia.

What is the probability of an accident which is not automobile?


Solution:

"n(S)=3500
\\n(X)=2800
\\ n(Y)=3500-2800=700
\\ P(Y)=\dfrac{n(Y)}{n(S)}=\dfrac{700}{3500}=\dfrac15"